# Large-scale online deanonymization with LLMs

### 18 Feb 2026 v1

@@ Abstract
We show that large language models can be used to perform at-scale deanonymization. With full Internet access, our agent can re-identify Hacker News users and Anthropic Interviewer participants at high precision, given pseudonymous online profiles and conversations alone, matching what would take hours for a dedicated human investigator. We then design attacks for the closed-world setting. Given two databases of pseudonymous individuals, each containing unstructured text written by or about that individual, we implement a scalable attack pipeline that uses LLMs to: (1) extract identity-relevant features, (2) search for candidate matches via semantic embeddings, and (3) reason over top candidates to verify matches and reduce false positives. Compared to prior deanonymization work (e.g., on the Netflix prize) that required structured data or manual feature engineering, our approach works directly on raw user content across arbitrary platforms. We construct three datasets with known ground-truth data to evaluate our attacks. The first links Hacker News to LinkedIn profiles, using cross-platform references that appear in the profiles. Our second dataset matches users across Reddit movie discussion communities; and the third splits a single user’s Reddit history in time to create two pseudonymous profiles to be matched. In each setting, LLM-based methods substantially outperform classical baselines, achieving up to 68% recall at 90% precision compared to near 0% for the best non-LLM method. Our results show that the practical obscurity protecting pseudonymous users online no longer holds and that threat models for online privacy need to be reconsidered.
@@ 1 Introduction
We demonstrate that large language models (LLMs) fundamentally change this calculus, enabling fully automated deanonymization attacks that operate on unstructured text at scale. Where previous approaches required predefined feature schemas, careful data alignment, and manual verification, LLMs can extract identity-relevant signals from arbitrary prose, efficiently search over millions of candidate profiles, and reason about whether two accounts belong to the same person. We show that the practical obscurity that has long protected pseudonymous users (the assumption that deanonymization, while theoretically possible, is too costly to execute broadly) no longer holds.
@@
We validate this thesis in three deanonymization settings: (1) deanonymizing an online account to its real identity; (2) linking an identity to an unknown pseudonymous account; and (3) linking pseudonymous accounts of the same person across different platforms or time periods. These settings capture distinct threat models (e.g., doxxing of an online account, a stalker targeting a victim, or an adversary consolidating a user’s activity across contexts) and pose different technical challenges.
@@
  • Extract: we ask LLMs to identify and structure relevant features from unstructured posts: demographics, writing style, interests, incidental disclosures, etc. Unlike prior deanonymization work that operated on predefined feature vectors (e.g., numerical movie ratings [20]), our approach captures the rich, heterogeneous signals present in natural language.
@@
Our findings have significant implications for online privacy. The average online user has long operated under an implicit threat model where they have assumed pseudonymity provides adequate protection because targeted deanonymization would require extensive effort. LLMs invalidate this assumption. They do not necessarily do so by exceeding human capability—the signals our models exploit are the same signals that a skilled investigator would recognize—but by reducing cost.
@@ 2 LLM agents can autonomously re-identify anonymous online profiles
To best highlight the paradigm shift that LLMs bring in online deanonymization, we begin our study by directly evaluating whether frontier LLMs can autonomously perform end-to-end deanonymization—from anonymous profile to real-world identity. This is a setting that has so far only been amenable to manual attacks, and included manually collecting hints and clues from online posts and performing strategic searches.
In order to evaluate whether LLMs can carry out this capability, we first need to build an evaluation set of online profiles along with a known ground truth identity. We do this by collecting non-anonymous online profiles and then explicitly removing any directly identifying information. Table˜2 summarizes the anonymization rules we use: we remove personal URLs, social media handles, GitHub repositories (which directly identify the owner) and retain institutions, demographics, interests, and names of connected people (which are too broad to identify the profile owner alone). The goal is to create a task that would be nontrivial for a motivated human with web search access. We then test whether LLM agents can solve this task by performing web searches. This artificial anonymization is only needed for evaluation (to establish ground truth); any real attack would target profiles that are already pseudonymous.
@@
    Summarize the user’s posts into a profile containing stated and inferred facts about demographics, career history, interests, and so on.
@@ 2.1 Results
We collect 338 Hacker News (HN) users who linked a LinkedIn profile in their publicly-visible HN bio, providing verified real-world identities as ground truth. We first summarize each user’s HN activity (comments and stories) into a structured profile. Then we create a search prompt and anonymize it (see Appendix˜A for details), and pass it to the agent. The agent correctly identifies 226 of 338 targets (67%) at 90% precision (95% CI: 86–93%; 25 incorrect identifications, 86 abstentions).
@@ 3 A framework for scalable deanonymization with LLMs
To better understand how LLMs deanonymize people, we need to explicitly state the different tasks that the LLM agent is solving. For this reason, we introduce a modular framework that decomposes deanonymization into distinct stages, each of which can be run with or without an LLM. This lets us scale to larger datasets, trace precision-recall curves, and ablate the contribution of LLMs at each step. Our framework is inspired by the seminal work of Narayanan and Shmatikov [20], whose attack serves as a “classical” baseline that we systematically augment with LLM components.
@@ 3.2 The ESRC deanonymization framework
    Calibrate: Finally, we need to calibrate the tradeoff between precision—how many of our predictions are correct—and recall—how many true matches we identify. The Netflix Prize attack does this by thresholding the similarity score of the top match to control the attack’s confidence (i.e., if the top match’s similarity is above a given threshold, the attacker guesses this match, and otherwise abstains). Similarly, we can use the embedding similarity score from the Search step, or an LLM-produced confidence score from the Reason step, or ratings computed by sorting matches via pairwise LLM comparisons. We use these confidence scores to trace precision-recall curves.
@@ 3.3 Experimental Setup
    Synthetically anonymized matches. We search for user profiles that explicitly link themselves to other platforms (e.g., a Hacker News user that posts their LinkedIn profile). Since these profiles trivially expose links, we carefully sanitize them by removing all identifying information.
@@ 4.2 Attack instantiation
We use the verification-stage confidence as the calibration score for precision-recall curves. For embedding-only and Narayanan baseline results, we use the gap between the top-2 candidates’ similarities—a large gap indicates the top candidate stands out clearly, making it more likely correct. The model outputs confidence via function calling, and we threshold it to trade off precision and recall, as shown in Figure˜3(a).
@@ 5.2 Attack instantiation
We use the verification-stage confidence as the calibration score for precision-recall curves; the selection-stage confidence is used only internally to pick the candidate passed to verification. The model outputs confidence via function calling, and we threshold it to trade off precision and recall. As a simpler alternative, we also evaluate using the gap between the top-2 embedding similarities as the confidence score—a large gap indicates the top candidate stands out clearly. As shown in Figure˜4(a), LLM confidence substantially outperforms this gap-based baseline.
@@ 6.2 Attack instantiation
To extract micro-data features, we use LLMs to filter and summarize the comments of each split-profile. We first apply a two-stage relevance filter: a heuristic pre-filter removes empty and deleted comments, very short responses, and pure URLs. Then, we prompt Gemini 3 Flash to label each of the remaining comments as relevant or generic. We discard generic comments and feed the remaining relevant ones to Gemini 3 Pro. The model generates a comma-separated list of the most important details, resulting in semi-structured micro-data. We discard both the query and candidate profile of users with zero comments after filtering (2 users) or if the LLM refuses to generate summaries due to inappropriate content (83 users). See Section˜G.2 for prompts and examples.
@@
Since similarity scores are poorly calibrated confidence measures, we sort all proposed query-candidate matches from most to least plausible using pairwise LLM comparisons. Concretely, we implement a Swiss-system tournament over all selected query-candidate pairs (Algorithm˜1), using the smaller GPT-5-mini model for efficiency. In each round, pairs of matches are compared head-to-head: the LLM sees two query-candidate pairs and judges which is a more plausible match. After each round, we update Bradley-Terry ratings [3] based on the comparison outcomes. We run 15 rounds with approximately 2,500 comparisons per round, and output matches sorted by their final rating.
@@ 7 Related work
Deanonymization attacks predate LLMs but were limited to structured data or high manual effort.
@@
Authorship attribution uses features such as function word distributions and syntactic structures [26]; recent work applies LLM embeddings to this task [29]. Our approach relies on semantic information—what users write about—rather than how they write; combining both is an interesting direction for future work.
The Extract step of our framework in Section˜3.2 follows a line of work of LLM inferring demographic data from online profiles. Staab et al. [24] show that LLMs can infer personal attributes such as location, occupation, and income from text with high accuracy, demonstrating privacy risks beyond training data memorization. Du et al. [8] extend this line of work with AutoProfiler, a system of four specialized LLM agents that collaboratively extract sensitive attributes from pseudonymous platform activity. They report 85-92% accuracy in attribute extraction, demonstrating that automated profiling can be deployed at web scale. However, their evaluation of actual deanonymization (linking extracted attributes to real identities) relies on manual LinkedIn searches and reports k-anonymity metrics rather than the high-precision matching we focus on. Conversely, Staab et al. [25] show that LLMs can be used to anonymize text while preserving utility, suggesting a potential defensive application of the same capabilities. Du et al. [9] survey emerging privacy risks from LLM deployment, distinguishing training-time data leakage from deployment-time threats such as automated profiling and social engineering.
Li [19] showed that agentic LLMs can deanonymize interview participants via web search, demonstrating that LLM agents make re-identification attacks low-effort: with a few natural-language prompts, off-the-shelf tools can search the web, cross-reference details, and propose matches. Their work on the Anthropic Interviewer dataset [15, 2] recovered 6 of 24 scientist identities by matching project descriptions to published papers, using task decomposition [17] to bypass safeguards. Nyffenegger et al. [22] evaluate LLM re-identification capabilities on court decisions, finding that despite high re-identification rates on Wikipedia, even the best LLMs struggled with anonymized legal documents. Our work extends this line of research by developing a systematic framework for LLM-based deanonymization and evaluating it across multiple platforms and attack settings.
@@ 8 Discussion
Deanonymization is one instance of LLMs acting as an “information microscope” that makes previously manual and expensive attacks scalable [14]. Our paper shows that LLMs democratize deanonymization; the asymmetry between attack cost and defense cost may force a fundamental reassessment of what can be considered private online.
@@
Classical anonymization frameworks such as $k$-anonymity [28] and differential privacy [10, 11] were designed for structured databases with explicit identifiers and assume attackers use direct matching or statistical queries. These frameworks do not account for the types of attacks we demonstrate; data releases should consider such threats when evaluating privacy risks. New frameworks could be developed with the threat of LLMs in mind.
@@
Our evaluation relies on ground truth datasets that may overestimate real-world success rates. Users who publicly link their accounts or share identifying information may still share more information than they would for truly pseudonymous accounts, even considering our anonymization step. However, measuring deanonymization performance requires ground truth, and we cannot verify matches for users who have not revealed their identities. Profile splitting has different limitations: users might behave differently across truly separate platforms than across different communities or across time on the same site. However, our methods work across a broad range of experimental setups, which suggests that they do generalize beyond any single evaluation setting.
@@ 9 Conclusion
Author Contributions
This research was directed by ETH Zurich, which conducted the primary experiments. Nicholas Carlini advised on the research.
@@ E.2 Reasoning effort comparison
There are two ways to use LLMs in the Calibrate step of our framework: (1) ask the LLM to output confidence scores directly alongside its match decision, or (2) sort all proposed matches by running pairwise comparisons. We use the first approach in Sections˜4 and 5, where the LLM outputs a confidence score via function calling during the Reason step, and we threshold this score to trace precision-recall curves. We use the second approach—Swiss-system sorting with Bradley-Terry ratings—in Section˜6, where it yields large gains over embedding similarity as a confidence measure.

### 25 Feb 2026 v2

@@ Abstract
We show that large language models can be used to perform at-scale deanonymization. With full Internet access, our agent can re-identify Hacker News users and Anthropic Interviewer participants at high precision, given pseudonymous online profiles and conversations alone, matching what would take hours for a dedicated human investigator. We then design attacks for the closed-world setting. Given two databases of pseudonymous individuals, each containing unstructured text written by or about that individual, we implement a scalable attack pipeline that uses LLMs to: (1) extract identity-relevant features, (2) search for candidate matches via semantic embeddings, and (3) reason over top candidates to verify matches and reduce false positives. Compared to classical deanonymization work (e.g., on the Netflix prize) that required structured data , our approach works directly on raw user content across arbitrary platforms. We construct three datasets with known ground-truth data to evaluate our attacks. The first links Hacker News to LinkedIn profiles, using cross-platform references that appear in the profiles. Our second dataset matches users across Reddit movie discussion communities; and the third splits a single user’s Reddit history in time to create two pseudonymous profiles to be matched. In each setting, LLM-based methods substantially outperform classical baselines, achieving up to 68% recall at 90% precision compared to near 0% for the best non-LLM method. Our results show that the practical obscurity protecting pseudonymous users online no longer holds and that threat models for online privacy need to be reconsidered.
@@ 1 Introduction
Deanonymization is a two-step process at heart, involving profiling an anonymous person from their posts, and then matching them to a known identity. It’s well-known that large language models (LLMs) can infer personal attributes from text on online forums [29, 8, 38]. Given this, it makes sense to ask: how good are LLMs at full end-to-end deanonymization, and is this a practical threat to pseudonymous accounts?
@@
We demonstrate that LLMs fundamentally change the picture, enabling fully automated deanonymization attacks that operate on unstructured text at scale. We show this by phrasing deanonymization as a matching problem and showing LLMs can perform all steps needed to match accounts: extract identity-relevant signals from arbitrary text, efficiently search over millions of candidate profiles, and reason about whether two accounts belong to the same person. We show that the practical obscurity that has long protected pseudonymous users (the assumption that deanonymization, while theoretically possible, is too costly to execute broadly) no longer holds.
@@
We validate this thesis in three deanonymization settings: (1) matching an online account to its real identity; (2) linking an identity to an unknown pseudonymous account; and (3) linking pseudonymous accounts of the same person across different platforms or time periods. These settings capture distinct threat models (e.g., doxxing of an online account, a stalker targeting a victim, or an adversary consolidating a user’s activity) and pose different technical challenges.
@@
  • Extract: as in prior work [29, 8], we ask LLMs to identify and structure relevant features from unstructured posts: demographics, writing style, interests, incidental disclosures, etc.
@@
Our findings have significant implications for online privacy. The average online user has long operated under an implicit threat model where they have assumed pseudonymity provides adequate protection because targeted deanonymization would require extensive effort. LLMs invalidate this assumption. As also observed by Staab et al. [29], they do not necessarily do so by exceeding human capability—the signals our models exploit are the same signals that a skilled investigator would recognize—but by reducing cost. Our experiments provide quantitative evidence for this in the deanonymization setting. We discuss our relationship to related work on online profiling and privacy in Section˜7; and address memorization concerns in Section˜8.
@@ 2 LLM agents can autonomously re-identify anonymous online profiles
To best highlight the paradigm shift that LLMs bring in online deanonymization, we begin our study by directly evaluating whether frontier LLMs can autonomously perform end-to-end deanonymization—from anonymous profile to real-world identity. This setting has so far required a significant manual component even if LLMs are used for attribute extraction [8].
In order to evaluate whether LLMs can carry out this capability, we first need to build an evaluation set of online profiles along with a known ground truth identity. We do this by collecting non-anonymous online profiles and then using an LLM to remove any directly identifying information. Table˜2 summarizes the anonymization rules we use: we remove personal URLs, social media handles, GitHub repositories (which directly identify the owner) and retain institutions, demographics, interests, and names of connected people (which are too broad to identify the profile owner alone). The goal is to create a task that would be nontrivial for a motivated human. We then test whether LLM agents can solve this task by performing web searches. This artificial anonymization is only needed for evaluation (to establish ground truth); any real attack would target profiles that are already pseudonymous.
@@
    Summarize the user’s posts into a profile containing both stated and inferred facts about demographics, career history, interests, and so on.
@@ 2.1 Results
We collect 338 Hacker News (HN) users who linked a LinkedIn profile in their publicly-visible HN bio, providing verified real-world identities as ground truth. We first summarize each user’s HN activity (comments and stories) into a structured profile. Then we create a search prompt and anonymize it (see Appendix˜A for details), and pass it to the agent. The agent correctly identifies 226 of 338 targets (67%) at 90% precision (95% CI: 86–93%; 25 incorrect identifications, 86 abstentions). We note that these edited profiles are much easier to identify than most pseudonymous accounts; we discuss this bias in Section˜3.3.
@@ 3 A framework for scalable deanonymization with LLMs
To better understand how LLMs deanonymize people, we need to explicitly state the different steps of successful deanonymization For this reason, we introduce a modular framework that decomposes deanonymization into distinct stages, each of which can be run with or without an LLM. While non-agentic, this approach lets us scale to larger datasets, trace precision-recall curves, and ablate the contribution of LLMs at each step. Our framework is inspired by the seminal work of Narayanan and Shmatikov [24], whose attack serves as a “classical” baseline that we systematically augment with LLM components.
@@ 3.2 The ESRC deanonymization framework
    Calibrate: Finally, we need to calibrate the tradeoff between precision (how many of our predictions are correct) and recall (how many true matches we identify). The Netflix Prize attack does this by thresholding the similarity score of the top match to control the attack’s confidence (i.e., if the top match’s similarity is above a given threshold, the attacker guesses this match, and otherwise abstains). Similarly, we can use the embedding similarity score from the Search step, or an LLM-produced confidence score from the Reason step, or ratings computed by sorting matches via pairwise LLM comparisons. We use these confidence scores to trace precision-recall curves.
@@ 3.3 Experimental Setup
    Synthetically anonymized matches. We search for user profiles that explicitly link themselves to other platforms (e.g., a Hacker News user that posts their LinkedIn profile). Since these profiles trivially expose links, we carefully sanitize them by removing all identifying information. Text sanitization is a well-studied problem [23, 30], though recent work shows that even sophisticated methods leave semantic signals that enable re-identification [38]; the point of our attacks is to exploit precisely the signals that mimic the traces that a genuinely anonymous user would leave in their profile.
@@ 4.2 Attack instantiation
We use the verification-stage confidence as the calibration score for precision-recall curves. For embedding-only and Narayanan baseline results, we use the gap between the top-2 candidates’ similarities—a large gap indicates the top candidate stands out clearly, making it more likely correct. The model outputs a numerical confidence score as a structured field in its function-call response, and we threshold it to trade off precision and recall, as shown in Figure˜3(a).
@@ 5.2 Attack instantiation
We use the verification-stage confidence as the calibration score for precision-recall curves; the selection-stage confidence is used only internally to pick the candidate passed to verification. The model outputs a numerical confidence score as a structured field in its function-call response, and we threshold it to trade off precision and recall. As a simpler alternative, we also evaluate using the gap between the top-2 embedding similarities as the confidence score—a large gap indicates the top candidate stands out clearly. As shown in Figure˜4(a), LLM confidence substantially outperforms this gap-based baseline.
@@ 6.2 Attack instantiation
To extract micro-data features, we use LLMs to filter and summarize the comments of each split-profile. We first apply a two-stage relevance filter: a heuristic pre-filter removes empty and deleted comments, very short responses, and pure URLs. Then, we prompt Gemini 3 Flash to label each of the remaining comments as relevant or generic. We discard generic comments and feed the remaining relevant ones to Gemini 3 Pro [14]. The model generates a comma-separated list of the most important details, resulting in semi-structured micro-data. We discard both the query and candidate profile of users with zero comments after filtering (2 users) or if the LLM refuses to generate summaries due to inappropriate content (83 users). See Section˜G.2 for prompts and examples.
@@
Since similarity scores are poorly calibrated confidence measures, we sort all proposed query-candidate matches from most to least plausible using pairwise LLM comparisons. Concretely, we implement a Swiss-system tournament over all selected query-candidate pairs (Algorithm˜1), using the smaller GPT-5-mini [28] model for efficiency. In each round, pairs of matches are compared head-to-head: the LLM sees two query-candidate pairs and judges which is a more plausible match. After each round, we update Bradley-Terry ratings [3] based on the comparison outcomes. We run 15 rounds with approximately 2,500 comparisons per round, and output matches sorted by their final rating.
@@ 7 Related work
Deanonymization attacks predate LLMs.
@@
The Extract step of our framework in Section˜3.2 follows a line of work of LLM inferring demographic data from online profiles. Staab et al. [29] show that LLMs can infer personal attributes such as location, occupation, and income from text with high accuracy, demonstrating privacy risks beyond training data memorization. Du et al. [8] extend this line of work with AutoProfiler, a system of four specialized LLM agents that collaboratively extract sensitive attributes from pseudonymous platform activity. They report 85-92% accuracy in attribute extraction, demonstrating that automated profiling can be deployed at web scale. However, their evaluation of actual deanonymization (linking extracted attributes to real identities) relies on manual LinkedIn searches and reports k-anonymity metrics rather than the high-precision matching we focus on. Conversely, Staab et al. [30] show that LLMs can be used to anonymize text while preserving utility, suggesting a potential defensive application of the same capabilities. On the other hand, Xin et al. [38] and Krčo et al. [19] show that even after removing explicit identifiers, contextual details in text still leak sensitive attributes, challenging the assumption that surface-level sanitization provides meaningful privacy.
More broadly, the privacy community has begun shifting focus from training-data-centric concerns (memorization, extraction) toward deployment-time threats such as automated profiling and inference [9, 22]. Our work provides large-scale empirical evidence for these deployment-time risks in the deanonymization setting.
Li [21] showed that agentic LLMs can deanonymize interview participants via web search, demonstrating that LLM agents make re-identification attacks low-effort: with a few natural-language prompts, off-the-shelf tools can search the web, cross-reference details, and propose matches. Their work on the Anthropic Interviewer dataset [16, 2] recovered 6 of 24 scientist identities by matching project descriptions to published papers, using task decomposition [18] to bypass safeguards. Apart from the mentioned AutoProfiler work on online platforms [8], several other recent works discuss re-identification in the context of redacted datasets. Nyffenegger et al. [26] evaluate LLM re-identification capabilities on court decisions, finding that despite high re-identification rates on Wikipedia, even the best LLMs struggled with anonymized legal documents. Xin et al. [38] propose a framework for evaluating re-identification risk in sanitized text datasets (medical records and chatbot conversations), finding that commercial PII removal leaves enough semantic signal for an adversary with auxiliary information to infer sensitive attributes. Our work extends this line of research by developing a systematic framework for LLM-based deanonymization and evaluating it across multiple platforms and attack settings.
@@
Authorship attribution uses features such as function word distributions and syntactic structures [31]; recent work applies LLM embeddings to this task [35]. Determining whether two texts were written by the same person has been studied extensively through the PAN shared task series [32], which evaluates pairwise verification on corpora of around 100 authors using stylometric features. Authorship attribution approaches explicitly treat topical similarity as a confound to be controlled for, whereas our method exploits semantic content as the primary signal. Our temporal-split Reddit setting (Section˜6) is related to authorship linking, but operates at larger scale (10,000 users) and relies on what users say rather than just how they say it. Combining stylometric and semantic approaches is an interesting direction for future work.
@@ 8 Discussion
Deanonymization is one instance of LLMs acting as an “information microscope” that makes previously manual and expensive attacks scalable [15]. Our paper shows that LLMs democratize deanonymization. Echoing concerns raised by prior work on LLM-based attribute inference and semantic privacy leaks [29, 38, 22], we argue that the asymmetry between attack cost and defense cost may force a fundamental reassessment of what can be considered private online. Our large-scale experiments provide quantitative evidence for these concerns in the deanonymization setting.
@@
Classical anonymization frameworks such as $k$-anonymity [34] and differential privacy [10, 11] were designed for structured databases with explicit identifiers and assume attackers use direct matching or statistical queries. These frameworks do not account for the types of attacks we demonstrate; data releases should consider such threats when evaluating privacy risks. LLM-based text anonymization [30] offers a more targeted defense for unstructured text, though even these methods leave residual semantic signals [38].
@@
Our evaluation relies on ground truth datasets that may overestimate real-world success rates. Users who publicly link their accounts or share identifying information may still share more information than they would for truly pseudonymous accounts, even considering our anonymization step. However, measuring deanonymization performance requires ground truth, and we cannot verify matches for users who have not revealed their identities. While there are synthetic benchmarks for attribute inference [39], it is not clear how to construct realistic synthetic datasets for deanonymization. Profile splitting has different limitations: users might behave differently across truly separate platforms than across different communities or across time on the same site. However, our methods work across a broad range of experimental setups, which suggests that they do generalize beyond any single evaluation setting.
@@ 9 Conclusion
Author Contributions and Acknowledgments
This research was directed by ETH Zurich, which conducted the primary experiments. Nicholas Carlini advised on the research. We thank Robin Staab and Mark Verö for extensive feedback.
@@ E.2 Reasoning effort comparison
There are two ways to use LLMs in the Calibrate step of our framework: (1) ask the LLM to output confidence scores directly alongside its match decision, or (2) sort all proposed matches by running pairwise comparisons. We use the first approach in Sections˜4 and 5, where the LLM outputs a structured confidence score during the Reason step, and we threshold this score to trace precision-recall curves. We use the second approach—Swiss-system sorting with Bradley-Terry ratings—in Section˜6, where it yields large gains over embedding similarity as a confidence measure.